# Formalization of Auslander–Buchsbaum–Serre criterion in Lean4

### v1

@@ 5.4.1. Auslander–Buchsbaum–Serre Criterion
With the results in  5.2 and  5.3, we can finally state the Auslander–Buchsbaum–Serre criterion as followsW.
@@
As noted in  2.5, global dimension of a ring $R$ equals the supremum of the global dimensions of the localizations $R_{\mathfrak{p}}$ as $\mathfrak{p}$ ranges over the prime ideals. Hence if $R$ has finite global dimension then each $R_{\mathfrak{p}}$ does as well. A useful corollary of the Auslander–Buchsbaum–Serre criterion is that the localization of a regular local ring at any prime is again regularW.
@@ 5.4.2. Regular Ring
We also developed basic theory of regular rings. The definition of a regular ring is simply that its localization at every prime is a regular local ringW.
@@
Using  2.5 and the fact that a regular local ring has global dimension equal to its Krull dimension, we obtain the analogous statement for regular ringsW.
@@
Since localization of a regular local ring at a prime is regular, regularity of a ring can be verified on maximal ideals onlyW.
@@
Finally, we prove that a regular local ring is a regular ringW.
lemma isRegularRing_of_isRegularLocalRing [IsRegularLocalRing R] : IsRegularRing R
@@ 5.4.3. Hilbert’s Syzygy Theorem
To prove Hilbert’s Syzygy Theorem (see [4, Corollary 18.1]), it remains to show that a polynomial ring over a regular ring is regularW.
@@
Because both regularity and the Cohen–Macaulay property are local conditions, the proof reduces similarly to the following local case: for a regular local ring $(R,\mathfrak{m},k)$ and a prime ideal $\mathfrak{p}$ of $R[X]$ with $\mathfrak{p}\cap R=\mathfrak{m}$, show that $R[X]_{\mathfrak{p}}$ is a regular local ringW.
lemma Polynomial.localization_at_comap_maximal_isRegularRing_isRegularRing
[IsRegularLocalRing R] (p : Ideal R[X]) [p.IsPrime] (max : p.comap C = maximalIdeal R) :
IsRegularLocalRing (Localization.AtPrime p)
One shows that $\mathfrak{p}$ can be generated by $0pt(\mathfrak{p})=\dim(R[X]_{\mathfrak{p}})$ elements, these generators then generate the maximal ideal of $R[X]_{\mathfrak{p}}$. If $\mathfrak{p}=\mathfrak{m}[X]$, it is generated by the minimal generators of $\mathfrak{m}$. Since $0pt(\mathfrak{m}[X])=0pt(\mathfrak{m})=\dim(R)$, the localization $R[X]_{\mathfrak{p}}$ is regular. If $\mathfrak{m}[X]\subsetneq p$, the image of $\mathfrak{p}$ in $k[X]$ a single polynomial $\overline{f}$. Lift $\overline{f}$ to $f\in R[X]$, then $\mathfrak{p}$ is generated by $f$ together with the minimal generators of $\mathfrak{m}$. As $0pt(\mathfrak{p})=0pt(\mathfrak{m})+1=\dim(R)+1$, the localization $R[X]_{\mathfrak{p}}$ is regular in this case as well. This completes the proof.
Thus we formalize Hilbert’s Syzygy theorem as followsW.
theorem Hilberts_Syzygy (k : Type u) [Field k] [Small.{v} k] (n : ℕ) :
globalDimension.{v} (MvPolynomial (Fin n) k) = n
@@ 6.1. Remarks About Implementation
In formalization of category theory in Lean4, one often needs to take universes into careful consideration. For example, in this project we must specify the universe in which the $\operatorname{Ext}$ functor lives. In our application we mainly use the instance for having $\operatorname{Ext}$ for categories with enough projectives W. For the category of $R$-modules in universe $v$ with $R$ in universe $u$, if we assume $R$ is $v$-small (denoted [Small.{v} R]), then we can obtain $\operatorname{Ext}$ in any universe greater than or equal to $v$. In our implementation, we only explicitly specify the universe level of $\operatorname{Ext}$ when it actually appears in the statement of a theorem; otherwise we hide it or fix a universe level. For example, for projectiveDimension and globalDimension, we use the standard universe following previous definitionsW. For the definition of depth, with $R$ in universe u and assuming Small.{v} R, in ModuleCat.{v} R we consider $\operatorname{Ext}$ in universe v, which is the smallest universe in which we obtain $\operatorname{Ext}$ using CategoryTheory.hasExt_of_enoughProjectives W. None of this matters much in practice, since there are already definitions for changing universes W, which simply pass through shifted hom, moreover, if the category is $R$-linear, then this is an $R$-linear equivalence and the additional structure is preserved.
A remaining nuisance is the universe of the modules on which we take $\operatorname{Ext}$. Since the commutation of $\operatorname{Ext}$ and Ulift is not fully formalized yet, some definitions still partially depend on a choice of universe. However, we know that $\operatorname{Ext}$ and Ulift do commute, so this will be resolved once that fact is formalized.
@@ 6.2. Related Works
As far as we are aware, in other theorem provers the concepts of derived categories and derived functors are often not formalized, let alone theories built on them such as $\operatorname{depth}$ and homological dimension. In Lean4, this work is also a first step toward applications of homological methods in commutative algebra.
For other related formalizations of commutative and homological algebra in Lean4, as mentioned in  2.3, there is a project that formalized that the Krull dimension of a polynomial ring over a Noetherian ring increases by exactly one (seeW). That proof proceeds by showing (a version of) [4, Theorem 19] : if $R\rightarrow S$ satisfies going down property, then the height of a prime ideal $P$ of $S$ equals the height of its preimage $\mathfrak{p}$ in $R$ plus the height of the image of $P$ in $S/\mathfrak{p}S$ (seeW). All the results about derived categories and $\operatorname{Ext}$ in mathlib that we use in this project originate from the Liquid Tensor Experiment development [2] and were ported to Lean4 [5].
@@ 6.3. Future Works
Building on the development of Cohen–Macaulay rings and regular local rings, there are several natural directions for future work.
Cohen–Macaulay Modules over Arbitrary Ring
For a non‑local ring $R$, one can define Cohen–Macaulay modules as follows: an $R$-module $M$ is Cohen–Macaulay if for every prime ideal $\mathfrak{p}$ of $R$, the localized module $M_{\mathfrak{p}}$ is Cohen–Macaulay over $R_{\mathfrak{p}}$.
@@
For a local homomorphism $R\to S$ with $R$ a regular local ring and $S$ a Cohen–Macaulay local ring, if $\dim S=\dim R+\dim(S/\mathfrak{m}_{R}S)$, then $R\to S$ is flat. This statement is known as miracle flatness (see [6, 00R4]). After the developments in this project, most of the required techniques are already formalized or are only a small step away. The remaining obstacle is formalizing a variant of the local criterion for flatness:
Lemma 6.1.
@@
We use the following standard notation for a ring $R$.
@@
Given our formalization of depth, we can state the criterion of normality as follows: for a Noetherian ring $R$, $R$ is normal iff it satisfies ($S_{2}$) and ($R_{1}$) (see [4, Theorem 39]). Formalizing this criterion would be an interesting project that will develop many results about normal domains along the way.

### v2

@@ 5.4.1. Auslander–Buchsbaum–Serre Criterion
With the results in Subsections  5.2 and  5.3, we can finally state and proof the Auslander–Buchsbaum–Serre criterion, following the original version in [1], as follows.W
Theorem 5.16.
A Noetherian local ring $R$ is regular if and only if $\operatorname{gldim}(R)$ is finite.
@@
As noted in Lemma  2.14, the global dimension of a ring $R$ equals the supremum of the global dimensions of the localizations $R_{\mathfrak{p}}$ as $\mathfrak{p}$ ranges over the prime ideals. Hence if $R$ has finite global dimension, then each $R_{\mathfrak{p}}$ does as well. This yields the following useful corollary.W
Corollary 5.17.
For regular local ring $R$ and $\mathfrak{p}$ a prime ideal of $R$, the localization $R_{\mathfrak{p}}$ is regular.
@@ 5.4.2. Regular Ring
We also developed the basic theory of regular rings.W
Definition 5.18.
A ring $R$ is a regular ring if $R_{\mathfrak{p}}$ is regular local ring for every prime ideal $\mathfrak{p}$.
@@
There is analogous statement of Theorem  5.4 for regular rings.W
Theorem 5.19.
For regular ring $R$, $\operatorname{gldim}(R)=\dim(R)$.
@@
Proof.
This is a direct consequence of Lemma  2.14 and Theorem  5.4. ∎
Since localization of a regular local ring at a prime is regular, regularity of a ring can be verified on maximal ideals only.W
Theorem 5.20.
A ring $R$ is regular if and only if $R_{\mathfrak{m}}$ is regular local ring for every maximal ideal $\mathfrak{m}$.
@@
Finally, we are able to prove that a regular local ring is a regular ring.W
@@ 5.4.3. Hilbert’s Syzygy Theorem
With all the previous results in this section we are able to prove a weak version of the Hilbert’s Syzygy Theorem. W
Theorem 5.21 (Hilbert’s Syzygy,[16, Corollary 18.1]).
For a field $k$, $\operatorname{gldim}(k[x_{1},\cdots,x_{n}])=n$.
theorem Hilberts_Syzygy (k : Type u) [Field k] [Small.{v} k] (n : ℕ) :
globalDimension.{v} (MvPolynomial (Fin n) k) = n
However due to the current setup, we still have some distance to the original theorem first stated in [11]. To prove the weak version above, it only remains to show that a polynomial ring over a regular ring is regular.W
Theorem 5.22.
For regular ring $R$, $R[X]$ is regular.
@@
Proof.
Because both regularity and the Cohen–Macaulay property are local conditions, the proof reduces similarly to the following local case: for a regular local ring $(R,\mathfrak{m},k)$ and a prime ideal $\mathfrak{p}$ of $R[X]$ with $\mathfrak{p}\cap R=\mathfrak{m}$, show that $R[X]_{\mathfrak{p}}$ is a regular local ring.W It suffices to show that $\mathfrak{p}$ can be generated by $0pt(\mathfrak{p})=\dim(R[X]_{\mathfrak{p}})$ elements, these generators then generate the maximal ideal of $R[X]_{\mathfrak{p}}$. If $\mathfrak{p}=\mathfrak{m}[X]$, it is generated by the minimal generators of $\mathfrak{m}$. Since $0pt(\mathfrak{m}[X])=0pt(\mathfrak{m})=\dim(R)$, the localization $R[X]_{\mathfrak{p}}$ is regular. If $\mathfrak{m}[X]\subsetneq p$, the image of $\mathfrak{p}$ in $k[X]$ is generated by a single polynomial $\overline{f}$. Lift $\overline{f}$ to $f\in R[X]$, then $\mathfrak{p}$ is generated by $f$ together with the minimal generators of $\mathfrak{m}$. As $0pt(\mathfrak{p})=0pt(\mathfrak{m})+1=\dim(R)+1$, the localization $R[X]_{\mathfrak{p}}$ is regular in this case as well. ∎
@@ 6.1. Remarks About Implementation
In the formalization of category theory in Lean4, one often needs to carefully consider universe levels. For example, in this project we must specify the universe in which the $\operatorname{Ext}$ functor lives. In our application, we mainly use the instance for having $\operatorname{Ext}$ for categories with enough projectives W. For the category of $R$-modules in universe $v$ with $R$ in universe $u$, if we assume $R$ is $v$-small (denoted [Small.{v} R]), then we can obtain $\operatorname{Ext}$ in any universe greater than or equal to $v$. In our implementation, we only explicitly specify the universe level of $\operatorname{Ext}$ when it actually appears in the statement of a theorem; otherwise we hide it or fix a universe level. For example, for projectiveDimension and globalDimension, we use the standard universe following previous definitionsW. For the definition of depth, with $R$ in universe u and assuming Small.{v} R, in ModuleCat.{v} R we consider $\operatorname{Ext}$ in universe v, which is the smallest universe in which we obtain $\operatorname{Ext}$ using CategoryTheory.hasExt_of_enoughProjectives W. None of this matters much in practice, since there are already definitions for changing universes W, which simply pass through shifted hom; moreover, if the category is $R$-linear, then this is an $R$-linear equivalence and the additional structure is preserved.
The remaining nuisance is the universe of the modules on which we take $\operatorname{Ext}$. Since the commutation of $\operatorname{Ext}$ and Ulift is not yet fully formalized yet, some definitions still partially depend on a choice of universe, such as moduleDepth, projectiveDimension and globalDimension. Currently we only now moduleDepth is invariant of universe in some good conditions by using the regular sequence characterization. However, as we know that $\operatorname{Ext}$ and Ulift do commute, it would be easy to prove that moduleDepth and projectiveDimension is invariant under arbitrary isomorphisms once that fact is formalized. As for globalDimension, from Lemma  2.13 we know that it suffices to consider the supremum over finitely generated modules. Notice that when assuming Small.{v} R and Small.{v’} R, there is a correspondence between finitely generated modules in ModuleCat.{v} R and ModuleCat.{v’} R. The globalDimension would be invariant when assuming the ring is small.
In our interaction with some categorical property and its corresponding versions stated in non-categorical language, for example Module.Projective, some related results had not previously taken universe levels into sufficient consideration. Fortunately, we fixed them when proceeding with our project.
@@ 6.2. Related Works
As far as we are aware, in other theorem provers the concepts of derived categories and derived functors are often not formalized, let alone theories built on them such as $\operatorname{depth}$ and homological dimension. In Lean4, this work also represents a first step toward applications of homological methods in commutative algebra.
Before this work, formalizations of commutative algebra in Lean4 have primarily been motivated by applications in algebraic geometry and algebraic number theory. For instance, the formalization of homogeneous ideals originated from the formalization of projective space [24]; developments on Dedekind domains and class groups were motivated by their roles in algebraic number theory [4]; and it is the formalization of local fields that furthered the formalization of valuation rings [7]. In contrast, the present work undertakes a systematic formalization of commutative algebra itself in combination with homological algebra, positioning it as a foundational body of theory upon which algebraic geometry and algebraic number theory can subsequently rely.
For formalizations of commutative and homological algebra in Lean4 directly related to this project, as mentioned in Subsection  2.3, there is a project that formalized that the Krull dimension of a polynomial ring over a Noetherian ring increases by exactly one W. That proof proceeds by showing (a version of) [16, Theorem 19]W. We would like to thank Andrew Yang here for his efforts in commutative algebra scattered throughout mathlib, which greatly benefited this work. All the results about derived categories and $\operatorname{Ext}$ in mathlib used in this project originate from the Liquid Tensor Experiment development [6] and were ported to Lean4 [19]. Even though we almost only used the definition of $\operatorname{Ext}$ and its long exact sequence, the work done at the level of derived categories has already laid a solid foundation for the subsequent formalization.
@@ 6.3. Future Works
Building on the development of Cohen–Macaulay rings and regular local rings, there are several natural directions for future works.
Cohen–Macaulay Modules over Arbitrary Rings
For a non-local ring $R$, one can define Cohen–Macaulay modules as follows: an $R$-module $M$ is Cohen–Macaulay if for every prime ideal $\mathfrak{p}$ of $R$, the localized module $M_{\mathfrak{p}}$ is Cohen–Macaulay over $R_{\mathfrak{p}}$.
@@
For a local homomorphism $R\to S$ with $R$ a regular local ring and $S$ a Cohen–Macaulay local ring, if $\dim S=\dim R+\dim(S/\mathfrak{m}_{R}S)$, then $R\to S$ is flat. This statement is known as miracle flatness (see [22, 00R4]). After the developments in this project, most of the required techniques are already formalized or are only a small step away. The remaining obstacle is formalizing a variant of the local criterion for flatness:
Lemma 6.1 ([22, 00ML]).
@@
We use the following standard notation for a ring $R$:
@@
Given our formalization of $\operatorname{depth}$, we can state the criterion of normality as follows: for a Noetherian ring $R$, $R$ is normal if and only if it satisfies ($S_{2}$) and ($R_{1}$) (see [16, Theorem 39]). Formalizing this criterion would be an interesting project that will develop many results about normal domains along the way.
Other Versions of Hilbert’s Syzygy Theorem
In Subsection  5.4.3, we proved that the global dimension of $k[x_{1},\dots,x_{n}]$ is $n$, which is only slightly stronger than the statement that every finitely generated module over $k[x_{1},\dots,x_{n}]$ has a projective resolution of length nn—the weakest version of Hilbert’s Syzygy Theorem. These projective resolutions can indeed be upgraded to free resolutions using the Quillen–Suslin theorem (see [18][23]) or other constructions specific to polynomial rings over fields (see [8, Corollary 19.8]). Additionally, there is a graded version of this theorem (see [8, Corollary 19.7]); however, with the current setup, it may not be feasible to formalize that version.